OpenStudy (anonymous):

Solve for x: |4x+12|=16

OpenStudy (anonymous):

Someone please help me!

OpenStudy (anonymous):

alright so we would first subtract 12 from 16 leaving four

OpenStudy (anonymous):

Answer choices A. X=7, x=-7 B. X=1, x=-1 C. X=1, x=-7 D. X=-1, x=7

OpenStudy (anonymous):

how many times would 4 go into 4? once. There for, the answer would be 1

OpenStudy (anonymous):

aishe

OpenStudy (anonymous):

Alright I understand

OpenStudy (anonymous):

So it's either b or c correct

OpenStudy (anonymous):

@RoxyKazzy

OpenStudy (lynfran):

|4x+12|=16 we need to get rid of the absolute value sign to do the we split the equation into 2 by (4x+12)=-16 and (4x+12)=16 we now solve each equations arriving at 2 solution

OpenStudy (lynfran):

for x

OpenStudy (texaschic101):

correct ^^....nice explaining Lyn

OpenStudy (anonymous):

Ok so my answer is c?

OpenStudy (lynfran):

so 4x+12=-16 we take 12 frm both sides we get 4x+12-12=-16-12 4x=-28 divide both sides by 4 we get x=-28/4 x=-7 now solve the other equation to get the other solution for x 4x+12=16 ??

OpenStudy (anonymous):

1?

OpenStudy (lynfran):

yes

OpenStudy (anonymous):

Ok thanks in a few mins can I ask u another question.

OpenStudy (lynfran):

sure

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Solve: -1/2x+1=-X+8

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

Answer choices A. 7 B. 1/2 C. 14 D. 18

OpenStudy (anonymous):

@texaschic101 @RoxyKazzy

OpenStudy (anonymous):

@LynFran please help your an amazing helper

Vocaloid (vocaloid):

-1/2 x + 1 = -x + 8 first add +x to each side of the equal sign

OpenStudy (anonymous):

Ok

Vocaloid (vocaloid):

what do you get after adding x to each side?

OpenStudy (lynfran):

yea like what vocaloid said ^^

OpenStudy (lynfran):

-1/2(x)+x+1=-x+x+8 1/2(x)+1=8 now subtract 1 frm both sides

OpenStudy (lynfran):

what do we get

OpenStudy (anonymous):

1/2x=7

OpenStudy (lynfran):

good now multiple both sides by 2

OpenStudy (lynfran):

we get??

OpenStudy (anonymous):

X= 14

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

Thank u so much!!!

OpenStudy (lynfran):

ok